OpenStudy (anonymous):

Solve. 2x2 + 3x = 20.

OpenStudy (anonymous):

ive gotten to (-3+/-root-151)/4 but I cant go anywhere from here... I cant find anything that -151 can simplify down to and my answers are not in radicals

OpenStudy (anonymous):

oh and its supposed to be 2x^2 sorry answer choices are:

OpenStudy (anonymous):

\[2x ^{2}+3x=20\] \[2x ^{2}+3x-20=0\] Then you do the quadratic equation: -3+/-√9-(-160)/4 -3+/-√(169)/4 -3+/-13/4 (-3+13)/4=10/4=5/2 (-3-13)/4=-4

OpenStudy (anonymous):

your answer would be x = 5/2 and x=-4 just in case, plug the answers back in to make sure they work

OpenStudy (anonymous):

so helpful, thank you!

OpenStudy (anonymous):

np, you just messed up the part inside the sqroot in the beginning that's all